OpenStudy (cookiemonster2412):

A train starts from rest and reaches a speed of 54 km/h in 10 sec. Find the acceleration

OpenStudy (anonymous):

acceleration is defined as a change in speed over a change in time. therefore, divide the change in speed by the time duration of the change a = (54 - 0) / (10) = 54/10 = 5.4 m/s^2